Hall Task List Spring17

Hall

• Change over water solenoids for the SS diode absorber trays. Water tank fitted – to connect.
• Nikhef Hall probes Melissa/Ajit – complete?
• Tracker vacuum. Ed stands
• Tracker turbo pumps. Craig, stands
• Luminosity monitor sensor 3 Craig/Ed - stands
• Luminosity monitor sensor 1 Craig/Ed - stands
• Diffuser iris 4 - re-test with field on 7/2/17. Stands
• ‘Splash guard’ for Tracker compressor plumbing in trench. Stands
• Tracker compressor change-over/testing. Stands

Code Updates C&M

• Ramp commands to SSU M coils – remedial action taken.
• Alarm handler/state machine. Moved to static config – expected complete end week 13th March.
• Water pump GUI password protected, #1891 – stands.
• Network errors messages – in progress

Electrical

Spectrometer solenoids
• Commissioning of Change over water solenoids for the SS diode absorber trays- T. Hartnett – complete, new tank installed, plumbing to be completed prior to re-test.
Focus Coil
• FC Data logger GUI to be tested and verified – ongoing
• PT100’s on FC in R9 to be fitted – ongoing 0 sensors will be glued
• FC power supply failure – repair with Sorensen
Hydrogen System
• Cryo-cooler for LH2 system moved from Vacuum control station 9 rack to H2 control room information to the control system is not available in new location, request to resolve. P. Hanlet to prove.
• Hydrogen level sensing commissioning – P. Warburton to advise
Drawings and records
• Check off power distribution records – J. Webb, A. Gallagher
• Records of testing to be made available - A. Gallagher
• Update all Mice drawings and issue to Mice Wiki page in PDF format – J. Webb, I. Mullacrane to advise
MICE Hall
• 24V supply for draw wire strain gauge required. Installation complete, commissioning and testing to be completed in hall. – A. Gallagher, J. Webb
• Magnet Interlock alarm is inoperative investigation required – DL engineer
Preventative maintenance
• Spectrometer solenoid Backup Helium Valve: Relay required to repeat valve state to the control system as status of valve is only known local to the valve. - I. Mullacrane
• Cable/Software check needed to prove which magnet Cold-heads are connected to which Cryo-compressors and that GUI has the correct compressor RS232 communications associated with the correct Cold-heads. - I. Mullacrane, J. Webb
• Review of FC cable numbers. - J. Webb
• Fit plinth to EMR rack. – A. Gallagher
• Temperature instrumentation upgrade for Decay Solenoid – I. Mullacrane
• RR2 to be updated in HevaComp – DL engineer
• Main 5th AC unit adjacent to mice substation reinstated, drawings require mods to show location and feeder details - J. Webb
• Stability Run for FC and Spectrometer Solenoid magnets – T. Hartnett, S. Griffiths

Detectors

  • New Target control PC (RR1) – in progress
  • Restart R78 target testing – in progress